# MAS-Bench: A Unified Benchmark for Shortcut-Augmented Hybrid Mobile GUI Agents

### v1

@@ 3.5 Systematic Hybrid Agent Evaluation
MAS-Bench is designed to evaluate an agent’s ability to autonomously discover and utilize shortcuts to improve task success, efficiency, and cost-effectiveness. This evaluation mandates that an agent possess sophisticated recognition and decision-making capabilities. Agents must not only identify available shortcuts based on the current task and environment but also critically assess their applicability and efficiency. Furthermore, the agent must make optimal decisions between GUI operations and shortcuts to minimize operational steps and time.
@@
We expect GUI agents not only to rely on predefined shortcuts within the knowledge base but also to possess the capability for shortcut generation. The quality of agent-generated shortcuts serves as a direct indicator of their generative capability. However, a systematic framework to develop and evaluate the effectiveness of these generation methods is currently lacking.
Therefore, MAS-Bench encourages and evaluates the agent’s ability to generate new shortcuts autonomously. As shown in Fig. 3, we design the following evaluation method. First, a self-generation shortcut agent explores the MAS-Bench environment and creates its shortcut knowledge base. To ensure fairness and eliminate interference from the intrinsic abilities of the agent under test, we use a unified baseline GUI agent as the evaluated agent and import these shortcut knowledge bases into the baseline agent for task execution. The baseline agent’s performance on the tasks directly reflects the quality of the imported knowledge base, thus enabling an evaluation of each agent’s shortcut generation capabilities.
@@ 3.6 Evaluation Metrics
We use the following key evaluation metrics to measure an agent’s performance on MAS-Bench comprehensively. We encourage agents to efficiently complete tasks by utilizing existing and self-generated shortcuts to achieve lower costs and greater efficiency.
Success Rate (SR).
Efficiency.
  • Mean Step Ratio (MSR): The ratio of the number of steps an agent takes to complete a task to the optimal number of steps required to complete the task. A lower ratio indicates better efficiency in task execution.
Cost and Resource Utilization.
  • Mean kTokens Cost (MToC): The average number of kTokens the LLM consumes during the task execution.
  • GUI to Shortcut Action Ratio (GSAR): This metric reflects an agent’s operational strategy preference by calculating the ratio of GUI to shortcut operations. A high ratio, when combined with low token consumption and short execution time, provides strong evidence that the agent utilizes shortcuts both efficiently and intelligently.
For automated evaluation, we adapt the methodology from SPA-Bench (Chen et al. 2024), embedding shortcut action descriptions to determine task success or failure. MAS-Bench provides a fair, reproducible, and in-depth assessment of a mobile agent’s comprehensive capabilities in complex GUI-Shortcut tasks through these multi-dimensional evaluation metrics. It specializes in agents’ potential for dynamic adaptation and innovative shortcut generation. By doing so, MAS-Bench offers valuable insights and direction for the future development of mobile agents.
4 MAS: Shortcut-Augmented Hybrid Agents
To evaluate the effectiveness of GUI-Shortcut hybrid operation, we introduce MAS-MobileAgent, which serves as a reference for GUI-shortcut Hybrid Agents. MAS-MobileAgent is based on MobileAgent-V2 (Wang et al. 2024b), which primarily relies on visual screenshots for UI understanding and automation. MAS-MobileAgent retrieves relevant shortcuts from a knowledge base and inputs them, along with perceptual information from screenshots, into the model to inform action decisions.
To test the generalization of our shortcut knowledge base and injection method further, we introduce MAS-T3A. Based on T3A (Rawles et al. 2024), this agent relies on a structured UI tree instead of visual information. By applying the same shortcut knowledge base to this agent with a different perceptual foundation, we aim to demonstrate the versatility of our approach, showing it can be seamlessly integrated into different agent architectures.
@@ 5 Experiments
Single-app Tasks (92 Tasks)
Human - 1.000 - - - -
T3A (Rawles et al. 2024) 0.511 1.056 137.641 346.382 0 0
M3A (Rawles et al. 2024) 0.565 1.064 192.775 155.281 0 0
MobileAgentV2 (Wang et al. 2024b) 0.446 1.058 1013.386 120.212 0 0
MobileAgent-E (Wang et al. 2025a) 0.359 0.818 459.574 88.772 0.378 0.081
MAS-T3A (Ours) 0.576 0.915 129.279 291.391 1.043 0.117
MAS-MobileAgent (Ours) 0.641 0.613 682.547 99.780 1.348 0.345
Cross-app Tasks (47 Tasks)
Human - 1.000 - - - -
T3A (Rawles et al. 2024) 0.340 1.087 257.122 625.970 0 0
M3A (Rawles et al. 2024) 0.383 1.262 411.145 288.833 0 0
MobileAgentV2 (Wang et al. 2024b) 0.170 1.247 2053.133 227.128 0 0
MobileAgent-E (Wang et al. 2025a) 0.064 0.934 469.109 85.859 2.250 0.177
MAS-T3A (Ours) 0.511 0.643 185.911 440.222 2.213 0.185
MAS-MobileAgent (Ours) 0.617 0.829 1441.586 189.836 3.128 0.320
Table 2: Performance comparison of our MAS agents and baseline methods on MAS-Bench with a predefined shortcuts knowledge base. All agents utilize the Gemini-2.5-Pro. $S$ and VH refer to the Screenshot and View Hierarchy (UI Tree) input modalities. MSRS is the Mean Step Ratio on Successful tasks, MET is the Mean Execution Time in seconds, and MToC is the Mean Token Cost in thousands (kTokens). The SSR for the predefined shortcuts is 1.0.
We conduct the experiments in the online, dynamic environment of MAS-Bench and evaluate the performance of GUI-shortcut hybrid mobile agent frameworks. We also employ an evaluation framework to measure the efficiency of agent-generated and hand-crafted predefined shortcuts.
@@ 5.1 Experiment Setup
Evaluation with Predefined Shortcut.
We evaluate the performance of GUI frameworks on the 139 tasks in MAS-Bench: 1) Frameworks with GUI-only action, including T3A (Rawles et al. 2024), M3A (Rawles et al. 2024), and MobileAgentV2 (Wang et al. 2024b). 2) self-generated shortcut frameworks, including MobileAgent-E (Wang et al. 2025a). 3) shortcut-augmented Hybrid GUI Agents (GUI-shortcut agents), including MAS-T3A and MAS-MobileAgent mentioned in Sec . 4 with predefined shortcut. We employ Gemini-2.5-Pro as the base model for all frameworks. The evaluation utilizes metrics from Sec . 3.6, with detailed configurations in Appendix A.
Evaluation of Shortcut Generation.
To assess the quality of shortcut knowledge bases generated by different strategies, we conduct a two-stage evaluation on a randomly selected test subset comprising 50% of the tasks in MAS-Bench (see Appendix C.2 for the task list). In the Shortcut Generation Stage, we construct several types of agent-generated shortcuts to form distinct knowledge bases. These are created using two methods: the first leverages MobileAgent-E to generate shortcuts ($S_{\text{MobileAgent-E}}$). The second, based on the historical execution trajectories of M3A, generates three types of shortcuts: task-level action replays ($S_{\text{Replay-Task}}$), subtask-level action replays ($S_{\text{Replay-Subtask}}$), and dynamic shortcuts ($S_{\text{Dynamic}}$ ) that require the model to perform real-time grounding, the construction of the agent-generated shortcut knowledge base is detailed in Sec.  3.3.
In the Quality Evaluation Stage, each knowledge base is mapped to a standardized action space and then integrated into the T3A baseline agent. This method treats the baseline agent’s performance as a direct reflection of the imported knowledge base’s quality, enabling an unbiased evaluation of the different generation strategies.
@@ 5.2 Effectiveness of GUI-Shortcut Operation
Based on our experimental results in Table 2, we identify three key findings:
Finding 1: Integrating a predefined shortcut knowledge base significantly enhances agent performance, thereby validating the effectiveness of the hybrid GUI-shortcut operation.
Experiments demonstrate that using just the predefined knowledge base of 88 shortcuts can significantly improve an agent’s task success rate and efficiency. For single-app tasks, compared to the baseline T3A, MAS-T3A improves the success rate (SR) from 51.1% to 57.6% and reduces the token cost from 346.382k to 291.391k. For agents that may terminate prematurely on long-horizon tasks due to errors, such as MobileAgent-E in the Table 2, the Mean Step Ratio on Successful tasks (MSRS) provides a more accurate assessment of the efficiency gains from shortcuts. MAS-MobileAgent achieves an MSRS of 0.613 on single-app tasks, which is significantly lower than the baseline version’s 1.058, indicating that its execution path is closer to the optimal solution. The SR and MS results for Different-level Tasks can be found in the Appendix B.2.
Finding 2: The effectiveness of predefined shortcuts is framework-agnostic and can benefit agents with different input modalities.
Experiments show that MAS-T3A and MAS-MobileAgent benefit from predefined shortcuts, regardless of whether they are based on UI Tree or screenshots. The reason lies in that the use of API, Deeplink, and RPA scripts is independent of the agent’s perception modality. These scripts directly send commands to the operating system or application, bypassing the need for the agent to parse the UI layout in real-time.
@@
Finding 3: The performance gain of shortcuts is more significant when the base model is weak.
For less capable models like Gemini-2.0-Flash, which often struggle with complex GUI navigation, shortcuts provide a crucial way to bypass failure-prone steps. This directly translates into a substantial increase in task success. For instance, on CROSS-app tasks, augmenting the Gemini-2.0-Flash agent with shortcuts boosts its success rate from 0% to 23.4% (see the Appendix B.3 for detailed experimental results). Shortcuts mainly contribute to efficiency improvements for strong models (e.g., Gemini-2.5-Pro). As shown in Fig. 5, the performance gain of using shortcuts is significantly greater for Gemini-2.0-Flash than for Gemini-2.5-Pro.
5.3 Result of Shortcut Generation
Our experiments include a baseline T3A framework without shortcuts and five shortcut variants: Predefined ($S_{\text{Predefined}}$), Replay-Task ($S_{\text{Replay-Task}}$), Replay-Subtask ($S_{\text{Replay-Subtask}}$), Dynamic ($S_{\text{Dynamic}}$), and MobileAgent-E ($S_{\text{MobileAgent-E}}$). Results demonstrate that shortcut quality significantly impacts both execution accuracy and efficiency.

### v2

@@ 3.5 Systematic Hybrid Agent Evaluation
MAS-Bench is designed to evaluate an agent’s ability to autonomously discover and utilize shortcuts to improve task success, efficiency, and cost-effectiveness. This evaluation mandates that an agent possess sophisticated recognition and decision-making capabilities. Agents must not only identify available shortcuts based on the current task and environment but also critically assess their applicability and efficiency. Furthermore, the agent must choose between GUI operations and shortcuts to reduce steps and time while avoiding irrelevant shortcut calls.
@@
We expect GUI agents not only to rely on predefined shortcuts within the knowledge base but also to possess the capability for shortcut generation. The quality of agent-generated shortcuts serves as a direct indicator of their generative capability. However, a systematic framework to develop and evaluate the effectiveness of these generation methods is currently lacking. Therefore, MAS-Bench encourages and evaluates the agent’s ability to generate new shortcuts autonomously. As shown in Fig. 4, we design the following evaluation method. First, a self-generation shortcut agent explores the MAS-Bench environment and creates its shortcut knowledge base. To ensure fairness and eliminate interference from the intrinsic abilities of the agent under test, we use a unified baseline GUI agent as the evaluated agent and import these shortcut knowledge bases into the baseline agent for task execution. The baseline agent’s performance on the tasks directly reflects the quality of the imported knowledge base, thus enabling an evaluation of each agent’s shortcut generation capabilities.
@@ 3.6 Evaluation Metrics
We employ 9 comprehensive metrics to evaluate agent performance across three dimensions: (1) Success: Success Rate (SR) measures task completion; (2) Efficiency: Mean Steps (MS), Mean Step Ratio (MSR), Mean Step Ratio on Successful tasks (MSRS), and Mean Execution Time (MET) assess operational efficiency; (3) Cost and Resource Utilization: Mean Token Cost (MToC), Mean Shortcut Call Count (MSC), Shortcut Success Rate (SSR), and Shortcut-to-GUI Ratio (S2GR) evaluate resource consumption and operational strategy. Table 8 in Appendix A.6 provides detailed definitions, units, and improvement directions for all metrics.
3.7 Evaluation Methodology
We introduce MAS-Bench-Eval as a scalable and reproducible evaluation protocol for MAS-Bench. Illustrated in Fig. 7, this automated pipeline adopts a two-stage Describe-and-Judge framework:
Stage 1: Action Semantics Extraction. An efficient MLLM processes the agent’s trajectory step-by-step, generating textual captions that describe the executed actions and corresponding UI state transitions. This stage transforms the raw visual stream into a structured semantic history.
Stage 2: Success Determination. A highly capable MLLM acts as the judge. It reasons over the task instruction, the semantic history derived from Stage 1, and the terminal screenshot to issue a binary success verdict.
This decoupled architecture effectively optimizes the trade-off between inference efficiency and evaluation capabilities. To validate its reliability, we compare MAS-Bench-Eval against 278 human-annotated trajectories and obtain high agreement with human judgments, with an F1 score of 96.3% and precision of 98.1%. We provide the full analysis in Appendix D.
4 MAS: Shortcut-Augmented Hybrid Mobile Agents
To evaluate the efficacy of the GUI-Shortcut hybrid operation, we introduce MAS-MobileAgent, which serves as a reference for GUI-shortcut Hybrid Agents. Built upon MobileAgent-V2 Wang et al. (2024a), which relies on visual perception for UI understanding, MAS-MobileAgent augments its decision-making process by retrieving relevant shortcuts from the knowledge base and integrating them with visual inputs.
To further assess the generalization of our shortcut injection mechanism, we introduce MAS-T3A. Derived from T3A Rawles et al. (2024), this agent operates on a distinct modality, utilizing structured UI trees rather than visual screenshots. By extending the hybrid operation to this text-based framework, we aim to verify its transferability to non-visual perceptual foundations.
Finally, to validate the effectiveness of hybrid operation within the Agent-as-a-Model paradigm, we apply the same shortcut injection strategy to multiple model families, including Qwen3-VL Bai et al. (2025a), GLM-4.5V Hong et al. (2025), ScaleCUA Liu et al. (2025c), and MAI-UI Zhou et al. (2025). Relevant shortcuts are incorporated directly into the system prompt without modifying the model weights. Collectively, these diverse implementations enable a rigorous and comprehensive assessment of our framework’s universality across mainstream agent architectures.
@@ 5 Experiments
[…]
Human - 12.11 1.000 - - - -
Agentic Workflow (Gemini-2.5-Pro)
M3A Rawles et al. (2024) 0.503 16.655 1.131 266.505 200.509 0 0
MobileAgent-E Wang et al. (2025d) 0.259 4.808 0.857 462.636 87.819 1.011 0.114
T3A Rawles et al. (2024) 0.453 15.755 1.067 178.005 440.919 0 0
   + MAS-T3A (Ours) 0.554↑22.3% 12.768↑19.0% 0.823↑22.9% 148.505↑16.6% 341.402↑23% 1.438 0.140
MobileAgentV2 Wang et al. (2024a) 0.352 19.252 1.122 1364.979 156.441 0 0
   + MAS-MobileAgent (Ours) 0.633↑79.8% 12.928↑32.8% 0.686↑38.9% 951.918↑30.3% 130.980↑16% 1.948 0.341
General-Purpose Models
Qwen2.5-VL-3B Bai et al. (2025b) 0.036 21.626 1.341 120.975 - 0 0
Qwen2.5-VL-7B Bai et al. (2025b) 0.022 21.835 1.498 168.832 - 0 0
Qwen3-VL-4B Bai et al. (2025a) 0.228 17.547 1.273 179.208 - 0 0
   + MAS-Qwen3-VL-4B (Ours) 0.237↑3.9% 20.403-16.3% 0.971↑23.7% 188.078-4.9% - 2.461 0.170
Qwen3-VL-8B Bai et al. (2025a) 0.259 16.576 1.183 179.377 - 0 0
   + MAS-Qwen3-VL-8B (Ours) 0.425↑64.1% 14.921↑10.0% 0.868↑26.6% 155.977↑13.0% - 1.309 0.110
Qwen3-VL-32B Bai et al. (2025a) 0.338 17.834 1.207 394.102 - 0 0
   + MAS-Qwen3-VL-32B (Ours) 0.446↑32.0% 15.913↑10.8% 0.848↑29.7% 358.282↑9.1% - 3.216 0.276
Qwen3-VL-235B Bai et al. (2025a) 0.417 16.604 1.111 185.201 - 0 0
   + MAS-Qwen3-VL-235B (Ours) 0.525↑25.9% 14.424↑13.1% 0.785↑29.3% 152.652↑17.6% - 2.784 0.233
GLM-4.5V Hong et al. (2025) 0.526 17.237 1.194 281.928 - 0 0
   + MAS-GLM-4.5V (Ours) 0.683↑29.8% 14.050↑18.5% 0.900↑24.6% 238.579↑15.4% - 1.051 0.097
Specialized GUI Models
UI-TARS-1.5-7B Seed (2025) 0.287 19.209 1.191 188.143 - 0 0
GUI-Owl-7B Ye et al. (2025) 0.295 15.568 1.239 168.148 - 0 0
ScaleCUA-7B Liu et al. (2025c) 0.108 18.194 1.202 123.018 - 0 0
   + MAS-ScaleCUA-7B (Ours) 0.115↑6.5% 19.446-6.9% 1.358-13.0% 141.535-15.1% - 0.007 0.000
ScaleCUA-32B Liu et al. (2025c) 0.231 19.209 1.286 141.654 - 0 0
   + MAS-ScaleCUA-32B (Ours) 0.216-6.5% 18.935↑1.4% 1.260↑2.0% 140.126↑1.1% - 0.000 0.000
MAI-UI-8B Zhou et al. (2025) 0.489 19.237 1.247 180.678 - 0 0
   + MAS-MAI-UI-8B (Ours) 0.583↑19.2% 18.065↑6.1% 1.179↑5.5% 169.463↑6.2% - 0.273 0.018
Table 2: Overall performance comparison on MAS-Bench with predefined shortcuts knowledge base (139 tasks). Results are weighted averages based on task distribution (92 single-app, 47 cross-app tasks). Bold and underlined values denote the best and second-best results within each block, respectively. Methods marked with “+” are shortcut-augmented variants of their corresponding baselines. $S$: Screenshot; VH: View Hierarchy; MS: Mean Steps; MSRS: Mean Step Ratio on Successful tasks; MET: Mean Execution Time (seconds); MToC: Mean Token Cost (in thousands); MSC: Mean Shortcut Call count; S2GR: Shortcut-to-GUI action Ratio. Detailed results in Appendix B.1.
We conduct the experiments in the online, dynamic environment of MAS-Bench and evaluate the performance of GUI-shortcut hybrid mobile GUI agents. We also employ an evaluation framework to measure the efficiency of agent-generated and predefined shortcuts.
@@ 5.1 Experiment Setup
We evaluate GUI agents on the 139 tasks in MAS-Bench, including three categories: 1) GUI-only agents, covering Agentic Workflows (e.g., T3A, M3A Rawles et al. (2024), and MobileAgentV2 Wang et al. (2024a)) and Agent-as-a-Model methods, including General-Purpose Models (e.g., Qwen3-VL Bai et al. (2025a) and GLM-4.5V Hong et al. (2025)) and Specialized GUI Models (e.g., UI-TARS-1.5 Seed (2025), GUI-Owl Ye et al. (2025), ScaleCUA Liu et al. (2025c), and MAI-UI Zhou et al. (2025)); 2) self-generated shortcut frameworks (MobileAgent-E Wang et al. (2025d)); 3) shortcut-augmented Hybrid Agents (GUI-shortcut agents), including MAS-T3A, MAS-MobileAgent, and MAS variants of Qwen3-VL, GLM-4.5V, ScaleCUA, and MAI-UI. For agentic workflows, we use Gemini-2.5-Pro as the base model. To examine the impact of predefined shortcuts across models with different capabilities, we also evaluate Gemini-2.0-Flash. All evaluations follow the metrics defined in Sec. 3.6 and Appendix A.
5.2 Evaluation Protocols
Predefined Shortcut Evaluation.
To evaluate GUI-shortcut hybrid operation, we augment each compatible GUI agent with the predefined shortcut knowledge base and compare it with its GUI-only counterpart under the same task set, environment, and evaluation metrics.
Shortcut Invocation Robustness.
To evaluate whether agents reliably select shortcuts, we annotate ground-truth shortcuts for each task and compare them with those invoked by agents. We report Mean Task Shortcut Recall (MTSR), Mean Shortcut Selection Precision (MSSP), MF1, and redundancy rate. To simulate long-tail cases where no suitable shortcut, we construct an interference setting on 23 cross-app tasks by removing ground-truth shortcuts and exposing task-irrelevant distractors. We report the changes in SR and MS, together with Mean Irrelevant Shortcut Calls (MISC). See Appendix B.4 for details.
Shortcut Generation Evaluation.
To assess the quality of shortcut knowledge bases generated by different strategies, we conduct a two-stage evaluation on a randomly selected test subset comprising 50% of the tasks in MAS-Bench. In the Shortcut Generation Stage, we construct several types of agent-generated shortcuts to form knowledge bases using two methods. One leverages MobileAgent-E to generate shortcuts ($S_{\text{MobileAgent-E}}$). The other, based on execution trajectories of M3A, generates three types of shortcuts: task-level action replays ($S_{\text{Replay-Task}}$), subtask-level action replays ($S_{\text{Replay-Subtask}}$), and dynamic shortcuts ($S_{\text{Dynamic}}$) that require the model to perform grounding. Details of the shortcut knowledge base are provided in Sec. 3.3. In the Quality Evaluation Stage, each knowledge base is mapped to a standardized action space and integrated into the T3A baseline agent. This method treats the baseline agent’s performance as a reflection of the imported knowledge base’s quality, enabling an unbiased, executor-controlled evaluation of the generation strategies.
@@ 5.3 Effectiveness of GUI-Shortcut Operation
Based on our experimental results in Table 2, we identify four key findings. Additional single-app and cross-app breakdowns for representative agents are provided in Appendix B.1.
Finding 1: Integrating a predefined shortcut knowledge base significantly enhances agent performance, validating the effectiveness of the hybrid GUI-shortcut operation.
Experiments show that using the predefined knowledge base of 88 shortcuts significantly improves task success rate and efficiency across multiple agent families. Compared to MobileAgentV2, MAS-MobileAgent improves the success rate (SR) from 35.2% to 63.3%, an 80% improvement, while reducing the token cost from 156.441k to 130.980k. The gain also holds for Agent-as-a-Model approaches: MAS-GLM-4.5V achieves the best overall SR of 68.3%, and MAS-Qwen3-VL-8B improves SR from 25.9% to 42.5%. Furthermore, the Mean Step Ratio on Successful tasks (MSRS) serves as a robust efficiency metric, particularly for mitigating the impact of premature termination in long-horizon tasks. MAS-MobileAgent achieves an MSRS of 0.686, which is significantly lower than the baseline version’s 1.122, indicating that its execution path is closer to the optimal solution. This also highlights the efficiency gap between current GUI-only agents and the human pure-GUI step baseline: strong GUI-only baselines remain above MSRS =1.0=1.0, while hybrid agents can reduce MSRS below 1.0 by bypassing GUI paths with shortcuts. The SR and MS results for Different-level Tasks can be found in the Appendix B.5.
Finding 2: The effectiveness of predefined shortcuts is framework-agnostic, benefiting agents with different input modalities.
Experiments show that both agentic workflows based on structured UI trees (MAS-T3A) and raw screenshots (MAS-MobileAgent), as well as Agent-as-a-Model variants (e.g., MAS-Qwen3-VL and MAS-GLM-4.5V), benefit from the same predefined shortcut knowledge base. This independence stems from the fact that shortcut execution is decoupled from the agent’s perception modality: shortcuts directly send commands to the operating system or application, reducing the need to parse the UI layout in real time. The results further show that General-Purpose Models adapt better to the hybrid action space, whereas specialized GUI models such as ScaleCUA do not automatically benefit without reliable shortcut selection.
@@
Finding 3: Shortcut gains depend on base model capability and shortcut-use reliability.
The Qwen3-VL series shows that shortcut gains are not monotonic with model scale: Qwen3-VL-8B achieves the largest relative SR gain (+64.1%), while Qwen3-VL-4B suffers an efficiency drop (+16.3% MS). This indicates that hybrid operation requires sufficient reasoning ability to select and invoke shortcuts reliably. At the same time, shortcuts can substantially help weaker but capable agents bypass failure-prone GUI steps; for example, augmenting Gemini-2.0-Flash boosts its cross-app SR from 0% to 23.4% (Appendix B.6).
Finding 4: MAS-Bench exposes and penalizes shortcut misuse.
Shortcut selection quality directly affects performance, as detailed in Appendix B.4. GLM-4.5V maintains low redundancy (4.8%) and reduces MS by 18.5%, whereas Qwen3-VL-4B has higher redundancy (33.5%) and increases MS by 16.3% (Table 11). In the no-relevant-shortcut interference setting, Qwen3-VL-4B suffers a 66.7% relative SR drop under irrelevant shortcuts, while GLM-4.5V drops only 7.7% (Table 12). These results show that MAS-Bench can evaluate and quantify agent robustness under noisy shortcut pools, not only performance when relevant shortcuts are available.
5.4 Results of Shortcut Generation
Our experiments include a baseline T3A framework without shortcuts and five distinct shortcut variants: Predefined ($S_{\text{Predefined}}$), Replay-Task ($S_{\text{Replay-Task}}$), Replay-Subtask ($S_{\text{Replay-Subtask}}$), Dynamic ($S_{\text{Dynamic}}$), and MobileAgent-E ($S_{\text{MobileAgent-E}}$). The predefined shortcut knowledge base serves as a reference upper bound for high-quality structural shortcuts, rather than a target that current automatic methods are expected to surpass. This setup allows us to compare shortcut generation paradigms under the same executor and quantify how close generated shortcuts are to reliable human-designed shortcuts.